Question 32 0 out of 1 points A 29.3 kg stone traveling at 6.97 m/s to the right, hits a stationary 17.6 kg ball. If the finaboth these questions have their answer given, i just need the steps to solve them. thanks.
Answer #1

using momentum conservation

m1 u1 + m2 u2 = m1 v1 + m2v2

29.3* 6.97 + 0 = 29.3* (- 3.56) + 17.6* v2

v2 = 17.53 m/s

====

33.

forces along vertical

T sin x = mg... (i)

foecws along horizontal

T cos x = m a.. . (ii)

dividing (ii) by (i)

tan x = g /a

a = 9.8 /tan 57.1

a = 6.34 m/s^2

=======

Comment before rate in case any doubt, will reply for sure.. goodluck